### v2

@@ 5.4 Individual Factors
5.5 In Depth Analysis
  • Decoder Limitations: The primary suspect for SAM’s subpar performance is the dependency of the image encoder on the prompt encoder and the mask decoder. To address this, we replaced the original mask decoder with a transformer decoder, as detailed in [15], modifying it for few-shot semantic segmentation following the inspiration from [41]. We introduced a constant prompt to the decoder, altering tokens to produce an $N$-class depth map rather than binary masks, and eliminated the IoU score output branch. We experimented with two mask decoder configurations, one with a Conv Transpose module and another where we upscaled the output through interpolation, while initializing remaining modules with pretrained weights. Despite these adjustments, this approach was outperformed by a simpler linear decoding method by up to 5% mIoU, likely due to the challenges of training a complex decoder in a Few-shot scenario.
  • Mask Distribution Bias: Another investigated issue was the mask distribution bias in COCO towards image centers, unlike the distribution in the SA-1B dataset used for SAM, highlighted in [15], where masks are more evenly distributed. We attempted to mitigate this bias with various preprocessing techniques (RandomCrop, CenterCrop, dividing each image into four smaller images); however, these modifications did not enhance our results.
6 Related Work
In this section, we discuss related work in the literature in few-shot classification and semantic segmentation.
6.1 Few-Shot Classification
Few-shot classification initially focused on in-domain tasks, with benchmarks involving dataset splitting into training, validation, and fake few-shot task generation for evaluation [30, 32, 37, 9]. The introduction of Meta-dataset [36] marked a shift towards the cross-domain setting, where the dataset used for model training differs from the tasks evaluated later, aligning with more realistic scenarios. More recently, the field has embraced the use of foundation models, adapting them to few-shot tasks. Notably, benchmarks such as those in [44] explore the application of CLIP [28] to 11 downstream tasks.
Methodologically, diverse approaches have emerged, ranging from hallucination-based [16, 11] to meta-learning [9, 25, 40, 24]. Straightforward methods have also demonstrated effectiveness [17, 35, 18]. Recent surveys, such as [21], suggest that optimal approaches often involve simple finetuning atop competitive pretrained models, augmented with an additional linear layer. This observation motivates the methods introduced in our paper.
6.2 Semantic Segmentation and Few-Shot Semantic Segmentation
Few-shot semantic segmentation is not a recent topic. Earlier works, such as [31] and [26], explored multiple benchmarks, including PASCAL-$5^{i}$ and COCO-$20^{i}$, reminiscent of in-domain benchmarks used in few-shot classification. Subsequent research ventured into using pretrained models on ImageNet [39], resembling the Meta-dataset approach in few-shot classification. Notably, most of these studies assumed binary labelling for the few-shots, emphasizing a single object of interest in the available shot images, considering all other objects as background.
In contrast, our approach considers that the provided few-shots are fully labelled, and we leverage pretrained foundation models, not limited to those trained solely on ImageNet. This approach aligns more closely with contemporary industrial requirements, focusing on the utilization of the most efficient available models, without specific explicit restrictions.
7 Conclusion
In conclusion, DINO V2 emerges as the top-performing backbone model, consistently outperforming other foundation models. While SAM exhibits strong overall results, it presents unexpected limitations on certain datasets. Surprisingly, our experiments underscore the competitive advantages of straightforward methods such as simple linear segmentation heads trained on the few given shots, challenging the necessity of complex procedures.
This work not only presents a comparison of techniques but also a practical benchmark. We hope this benchmark will contribute to improving few-shot semantic segmentation methods and will provide a solid ground for comparison.

### v3

@@ A.5 More Individual Factors
A.6 Model Comparison Summary
Table 10 summarizes the main components of the different models used in our study, we discuss the importance of each component (Architecture, Size, Pre-training Dataset, Pre-training Method) in 6.3.
Table 10: Summary of the main model elements.
Architecture Size Dataset Pre-training Method
SAM ViT 89M SA-1B MAE + Prompt Segmentation
DINOv2 ViT 86M 142M image dataset DINO (SSL with Knowledge Distillation)
iBOT ViT 85M ImageNet-1K Masked image modeling + knowledge distillation
CLIP ViT 86M 400M of (image,text) pairs image, text embedding alignement
MAE ViT 86M ImageNet-1K Masking + Image reconstruction
FCN-ResNet50 CNN 23M Subset of COCO Semantic Segmentation
SegFormer ViT 81M ADE20K Semantic Segmentation
A.7 Dataset Comparison Summary
Table 11 and Figure 4 provide an overview of the three datasets considered in this study. These datasets exhibit varying levels of class diversity. Notably, in the PPD dataset, segmentation masks show significant variation across images, contributing to a higher standard deviation in mIoU. This variability poses a challenge in one-shot learning scenarios, where models must generalize effectively from only two reference images.
Table 11: Overview of the datasets used in our benchmark.
Dataset Number of Classes Resolution Size of Training Set
Cityscapes 19 1024×2048 2,975
COCO 80 (400–640)x(400–640) 118k
PPD 2 500×500 783
Cityscapes
COCO
PPD
Figure 4: Examples of images and corresponding segmentation masks from Cityscapes, COCO, and PPD datasets.
A.8 Impact on the Number of Shots
We can see in Fig. 5 that for SVF, the different models scale similarly and that models are still outperformed by DINOv2 even when more shots are available. We can see in Fig. 6 that for DINOv2, the gap between the models tends to be smaller the more we increase the number of shots. Also, full fine-tuning scales better than other methods because we are less prone to overfitting, leading to the conclusion that the is it better to choose the full fine-tuning method when more shots are available. Yet it is not crucial since the full fine tuning is time and memory expensive compared to the other methods, and a simple linear probing can give good performance with a minor drop in performance.
1251020304050607080Number of shotsmIoUDINOv2SAMCLIPMAEResNetiBOTSegFormer
Figure 5: Impact of the number of shots on mIoU for the SVF method.
1251050607080Number of shotsmIoULinearMultilayerSVFLoRAFine-tuning
Figure 6: Impact of the number of shots on mIoU for DINOv2.
A.9 COVID-19 CT Dataset
We conducted additional experiments on a fourth dataset, a medical dataset for COVID-19 CT image segmentation (https://www.kaggle.com/c/covid-segmentation), which includes 4 classes. The 1-shot results (mIoU), presented below, demonstrate a significant performance drop for Vision Foundation Models (VFMs) trained predominantly with SSL methods on natural images when applied to medical images. This gap in performance, caused by the substantial distributional differences between natural and medical images, highlights the advantages of current semantic segmentation models, such as ResNet-50 and SegFormer, in handling tasks with large domain shifts.
Table 12: mIoU averaged over our three considered datasets in 1-shot semantic segmentation for COVID-19 CT Dataset. Bold numbers correspond to the best performance across extractors and underline numbers to the best performance across adaptation methods.
Model Linear Multilayer SVF LoRA Finetune
SAM 23.20±02.37 45.69±05.40 32.90±03.21 39.28±05.25 24.49±01.39
DINOv2 34.74±02.52 40.18±02.35 30.11±07.36 34.05±00.76 32.01±04.82
CLIP 34.95±01.45 35.50±01.17 36.70±02.40 34.71±00.83 26.98±01.61
MAE 23.37±01.53 24.10±01.75 23.48±01.83 23.92±01.75 21.04±01.22
FCN-ResNet50 44.01±01.27 54.03±01.14 56.97±01.05 54.71±01.41 57.35±02.14
IBOT 23.51±03.19 24.54±03.02 21.87±00.46 40.28±03.79 21.49±01.63
SEGFORMER 44.14±00.82 50.30±01.20 49.35±01.02 47.51±01.49 52.25±00.99
A.10 Limitations
While our benchmark introduces novel insights into the adaptation of Vision Foundation Models (VFMs) for few-shot semantic segmentation, several limitations warrant further investigation:
  • Dataset Representation: Our benchmark relies on a limited number of datasets, primarily from urban scenes (Cityscapes), natural images (COCO), and plant phenotyping (PPD). Although diverse, this selection does not encompass other important domains, such as satellite imagery or underwater scenes, where few-shot segmentation could be valuable.
  • Model Adaptation Scope: The study focuses on a specific set of pretrained VFMs and adaptation methods. While comprehensive, this excludes potential alternative models, such as specialized medical imaging networks or models pretrained on multimodal datasets, which could perform better in domain-specific settings.
  • Domain Gap: The significant drop in performance observed for medical images (COVID-19 CT dataset) highlights that VFMs trained on natural images are not inherently robust to large domain shifts. Addressing this limitation may require dedicated pretraining or domain adaptation strategies, which were outside the scope of this study.
  • Energy Efficiency: The absence of an energy efficiency analysis is a limitation of our work. However, it is evident that methods like LoRA and SVF are considerably more efficient than full fine-tuning, as they require fewer computational resources and less processing time while maintaining strong performance. A more detailed study on the energy consumption of different adaptation techniques would be valuable for future research.
These limitations present opportunities for future research to enhance the applicability and robustness of VFMs in few-shot semantic segmentation, particularly in challenging, real-world scenarios.
A.11 Negative Societal Impact
VFMs are trained on carefully curated large-scale datasets designed to exclude harmful content and private information, mitigating privacy risks a priori. These datasets are generally less biased than traditional datasets, helping reduce dataset-induced biases in model predictions. However, potential biases and harmful content may still arise, necessitating continuous efforts in debiasing and content filtering. Additionally, while the benchmark we developed has minimal potential for direct malicious applications, the models used in our research could still reflect biases from their pretraining or adaptation datasets. Furthermore, by streamlining semantic segmentation and enabling effective model training with limited labeled data, this approach could inadvertently lower the barrier for applications in surveillance or other privacy-sensitive domains. While this accessibility fosters innovation and practical deployment, it also highlights the importance of responsible use and potential regulatory considerations in sensitive contexts.
A.12 Object Size Effect on mIOU
We analyzed the relationship between object size and segmentation performance by plotting the mIoU for the Car class against its pixel area (see Figure 7). Our results show that larger cars tend to achieve higher mIoU scores, while smaller cars exhibit lower mIoU and higher variance.
This trend may not solely reflect better model performance but could be due to a geometric bias: larger objects have more inner pixels, which are easier to segment, whereas smaller objects have a higher proportion of boundary pixels, which are harder to segment accurately.
These findings suggest that object size influences segmentation performance, but the apparent advantage for larger objects may stem from intrinsic biases rather than true model generalization.
[…]
Figure 7: Impact of Car Area on mIoU for the ’Car’ class using DINOv2.